The formula for Converting Fahrenheit temperature to Celsius is C=5/9(F - 32)
If the Celsius temperature on a certain Winter Day Minneapolis is neverless than -25 degrees how would you describe the corresponding Fahrenheit temperatures? (please help in detailed steps thank you :) )

Answer by stanbon(75887)

The formula for Converting Fahrenheit temperature to Celsius is C=5/9(F - 32)
If the Celsius temperature on a certain Winter Day Minneapolis is never less than -25 degrees how would you describe the corresponding Fahrenheit temperatures?
-----------------------
INEQUALITY:
(5/9)(F-32)>=-25
Multiply both sides by (9/5) to get:
F-32 >= -25(9/5)
F-32 >= -45
Add 32 to both sides to get:
F >= -13
===============
Cheers,
Stan H.